EVA Report – January 03rd

  


EVA #: 12

Autor: Luz Maria Agudelo Urrego

Purpose of EVA: weather observation; Radio measurements; EVA and EMU evaluation project; Reliability and maintenance

Start time: 10:18

End time: 12:25

Narrative: Shefali, Ben and Luz Ma, drove Curiosity and Spirit along Cow Dung Rd, and stop before the original destination, Brahe, to assure the charge of the rovers Curiosity and Spirit was enough for the way back to the Habitat. They took weather observation, radio measurements, performed a reliability scenario, and explored the canyon. An engineering EVA was performed at the end of the excursion.

Destination: Along Cow Dung Rd

Coordinates (use UTM NAD27 CONUS): 12S 517950, 4253860

Luz Ma
ENG
Crew 218

Science report – January 03rd

Crew 218 Science Report 03-JAN-20
Crew Science Officer: Dr. Jonathan R. Buzan

1. Decision Making in support of autonomy for crew EVAs:

No EVA: nothing to report.

2. Mars surface weather:

Measurements were collected during today’s EVA.

3. Subsurface structure on Mars:

No EVA: nothing to report.

4. Detecting radio signal strength:

Field measurements were collected during today’s EVA.

5. EVA workload analysis:

Survey data was collected for today’s EVAs.

6. EMU (Extravehicular Mobility Unit) ergonomic assessment:

Survey data was collected for today’s EVAs.

7. Environmental Stresses over MDRS habitat and Crew Members and projection over Martian Terrain:

Interior measurements were expanded to 2nd floor hab. Plans to take measurements at night with state room door open.

8. Messier and other space objects for outreach:

Cloudy: nothing to report.

9. Reliability and maintenance:

Simulated malfunction of ENG. LuzMa’s radio.

10. Medical readings in preparation for future crew-wide project:

No EVA: nothing to report.

11. Collection of clay, shale, and hematite samples:

No EVA: nothing to report.

12. Media and outreach:

Nothing to report.

Glassware check out: None

A reminder to all crewmembers: There’s a $300 fine for using any glassware material without MDRS Mission support permission.

GreenHab Report – January 03rd

  


Crew 218 GreenHab Report 03-JAN-20

Crew GreenHab Officer: Dr. Jonathan R. Buzan

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 21°C; 19%

11:15AM

Floor Unit: 15°C

Electronic: 23.6°C

humidity 18%

3:00PM

Floor Unit: 15°C

Electronic: 20.5°C

humidity 20%

6:40PM

Floor Unit: 14°C

Electronic: 19.9°C

humidity 20%

Max: 25.1°C; 22%

Min: 16.9°C; 17%

Hours of supplemental light: Light system 6:00PM-11:00AM.

Daily water usage for crops: 6.4Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – ~74 Gal.

Time(s) of watering for crops:

6:40PM

SEASONAL PLANTING HISTORY:

Change to crops: None.

OTHER NOTES:

1st sprouts: N/A

—Gnat Spray used on all plants.

—Snow peas have dried out. Soil was moist. Not sure how to proceed.

Harvest: N/A

Sol Summary – January 03rd

  

Sol: 12

Summary Title: Mars Stars

Author's name: Pat Pesa

Mission Status: Sad that our mission is coming to a close.

Sol Activity Summary: Today we had our last EVA ( :( ) and then started the long process of packing and cleaning for the next crew coming tomorrow.

Look Ahead Plan: Tomorrow we look forward to finishing cleaning and welcoming Crew 219!

Anomalies in work: None

Weather: Foggy and cloudy, around 30 degrees

Crew physical status: Healthy

EVA: EVA #12 this morning was our last EVA of the mission. They went north on Cow Dung Road for continued radio signal and weather observations.

Reports to be filed: sol summary, commander report, operations report, greenhab report, journalist report, EVA report, science report

Support Requested: none

Pat Pesa

Geologist, MDRS 218

Commander Report – January 03rd

  


Crew 218 Commander Report – 03-01-2020

Sol 12 – Good things come to an end? Good things go into better things!

As usual, time during a mission at MDRS flies. So, we are already on our way back to Earth, ready to open our hatch and come out of simulation tomorrow. It has been a fantastic ride, with different challenges than in my previous missions, mostly weather-related, and a lot of fun and good work. The feelings are mixed: we are a little tired and probably ready to come back to our friends and families, though we know that the first impact with crowds will be surprising. There is obviously some sadness, as it is for every good thing that comes to an end. But my feeling is that the kind of good things that we do at MDRS do not end here, but they evolve into better things. This is true not only for those who, like me, are veterans of MDRS. Everybody who is part of a crew at MDRS goes back home with more inspiration to be spread around, with a better understanding of the challenges of space exploration and of oneself, and with the pride of having given a personal contribution to the great endeavor we all dream of: expanding our horizons, both personal and as humankind, to be able to leave Earth on the route of exploration and knowledge and adventure.

My gratitude to the many people who supported our mission!

Cesare Guariniello, Commander

Operations Report – January 03rd

  

Journalist Report
Ben Durkee, Crew 218 Journalist

Sol 12

Today I was informed that the GreenHab isn't soundproof. The past few sols I've been going in there and belting out some Billy Joel while photographing and hanging out with the plants. I'm not quite homesick yet, but I would say I'm becoming a bit pianosick, if that makes any sense. Keeping my pipes warm helps keep me from going space crazy, I think - maybe it helps the plants grow too. Only now do I find out that the rest of the crew heard everything. I've been betrayed by the only other people on the planet. I'll show them. Once I become the first Martian Disney princess, they'll see. They'll all see.

In other, somewhat bittersweet news, today was our last EVA! EVA 12 on Sol 12; it was very exciting, featuring both ends of the expeditionary spectrum. We gathered great data and explored a beautiful part of the landscape, but we also encountered considerably more technical difficulties than on any of the other EVAs. LuzMa, Shefali, and I set off after sunrise and took the rovers North. We continued our trajectory until one of the rovers was just above the "return home" threshold and we had found the spot for our surveys. We each gathered the last data we would get for our Mars research projects and returned to the rovers. In the midst of our data collection, we hit a few road bumps.

The EMU suits are perfect for keeping us alive and (somewhat) mobile on the Martian surface. However, they have one Achilles' heel: the fog problem. Because of the humidity of our breath, the transparent surface of the helmet begins to fog up progressively, even with air continually flowing into the helmet. This problem can be mostly circumvented by wearing a bandana or scarf which will divert your breath away from the view-bubble, but even then some condensation escapes. Additionally, this strategy is not foolproof. If your face mask falls down during any point in the EVA, it's game over. You can't reach into the helmet to adjust or re-secure it, and all my attempts to recover it by mouth have been unsuccessful (and presumably very entertaining).

We found that something about the weather conditions today made the helmets particularly ripe for fogging. Compound that with the rocky drive which shook off all of our masks, and you've got a recipe for three blind astronauts. While we walked around gathering our data, the exertion of trudging through snow in the EMUs had our helmets looking like mystical crystal balls before long.

When we regrouped at the rovers, we found that we still had quite a bit of time remaining in our EVA window. We rested for a few minutes and regulated our breathing in an attempt to mitigate the fogging as much as possible. It worked, temporarily. It was at this point where we simulated a radio malfunction for Shefali's human factors research. We ran through communication loss protocol, in the midst of which we discovered a real equipment fault - a wardrobe malfunction.

I don't know if you've ever tried to tie someone's shoes with gloves on, but it's impossible. Now try with only 50% of your normal vision and 30% of your standard range of motion. Through the magic of Mars, one of the first things we learn as children becomes the hardest task imaginable. After trying profusely to tie a traditional bow, I gave up and just tied three overhand knots (the normal one that is the basis of a bow knot). Yes, I used some of our very limited bandwidth to google what that knot is called. Maybe I should've become a Boy Scout, after all.

We got everything sorted out and decided to use our remaining EVA time to explore the surrounding area. We marched through the harsh marsh of Martian snow and discovered a beautiful grotto hidden away in the rocks. It had a frozen stream flowing into it which terminated under an enchanting overhang of red rock. Beautiful orange icicles cascaded down from the overhang, appearing to be a set of alien vines. Below those gathered some lovely frozen foliage and picturesque rocks, all seeking shelter from the sun. I can only imagine what that nook what look like on warmer days.

I suggested that we dub it Bean Grotto, in honor of the bean detritus and bean puns that have been plaguing our crew as of late. The comms chatter blew up with groans and firm "No!"s. I didn't hear anyone suggest a different name, though...

In our excursion to the desert cove, our helmets continued fogging up with haze that just would not disappear, no matter how hard we tried. We eventually had to return to the rovers to try and solve the issue, or worst case allow us enough time to drive home in spite of the fog problem. We attempted to regulate our breathing and to wipe the glass with our faces, but to no avail. As our last resort, we had to break simulation to clear our visors of obstruction. It was unfortunate, but safety comes first and you don't mess around with vehicles in the unforgiving Martian desert.

After a cautious, rugged drive home and a few pauses for the fog to clear once more, we all made it back to the Hab safely. We regaled the crew with the tale of our heroics and spent a short while to catch our breath. Then came time to begin Habitat clean-up.

It's wistful, having to gather up our belongings from the giant terrarium that we've just grown accustomed to living in. This place had just become our home, and now we have to return it to its original state so we can pass the torch to our successors. In spite of the melancholy, I'm super stoked to meet the astronauts who will soon inherit the station. Their pod lands tomorrow afternoon, and we are ready to welcome them with open arms. It will be nice to see some fresh faces after two weeks! And then bombard them with inside jokes and the quirks of Martian living that we've discovered.

After some more spring cleaning and then a dinner assembled of the remaining food scraps, we will retire for our last night alone in the Hab. It's been one hell of a ride, that's for sure. These two weeks have been unbelievably dense with learning, bonding, and growing as individuals. The Ben that arrived on Mars and the Ben that will blast off on Sunday are two drastically different people. This has been an incredibly valuable experience, and I don't think I'll know what to do with myself once I'm back on Earth. I guess we'll cross that bridge when we come to it.

Now, it's time for me to grab the broom and head back to the GreenHab for a little more quality time with my plant pals. At least they appreciate my singing.

	

Operations Report – January 03rd

  


Crew 218 Operations Report 3-Jan-20

SOL: 12

Name of person filing report: Luz Maria Agudelo Urrego

Non-nominal systems: NA

Generator: run

Hours run: 14

From what time last night: 1730

To what time this morning: 0730

List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before the generator is run at night: 62%

Diesel Reading – 73%

Station Propane Reading – 60%

Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 50

Water Meter (units): 0147279.8

Water (static tank) (gal): 225

Static to Loft Pump used – No

Water in Green Hab (gal): 74

Water in Science Dome (gal): 0

Toilet tank emptied: No

Deimos rover used: Still in the workshop

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: Assigned to director

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: EVA 12

Hours: 122.7

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging): 43%

Currently charging: Yes

Opportunity rover used: Still in the workshop

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: EVA 12

Hours: 128.1

Beginning charge: 100%

Ending charge: (On return from EVA, before charging): 42%

Currently charging: Yes

Notes on rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use: N/A

Oil Added? No

# Hours the ATVs were used today: 0

Notes on ATVs: N/A

HabCar used, and why, where? No

CrewCar used, and why, where? No

General notes and comments: Nominal

Summary of the internet: Nominal

Summary of suits and radios: We followed the instruction from Scott Davis and fixed the Suit N°6 on the two-piece space simulator suits. Suit N°6 was used today during EVA 12, and it works perfectly.

Two of the Helmets need to be repaired. During EVA 11 preparation, we found out a helmet was missing a connector, and the red helmet had the back broken before we arrived. We put a note on the helmets for distinction.

Summary of Hab operations: The pipe from the static tank through the loft tank is still frozen. The crew filled out the Loft Tank by carrying water buckets from the Static Tank.

Summary of Science Dome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: Nominal

Summary of health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: NA

Luz Ma
ENG
Crew 218